How do I solve for n? C(n, 5) = C(n, 7)
Found 2 solutions by solver91311, richard1234:
Answer by solver91311(24713)



If you consider all of the possibilities for various combinations of things taken at a time and listed them in desceding order of , you would get a list that looks like:



















(Since , these two list items must be equidistant from the middle item in the list.)


















But since the s must all be integers, there can only be one item in the list between 7 and 5, namely



hence



So




John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by richard1234(7193)

Use the identity that C(n, k) = C(n, n-k). Here, we can let k = 5, n - k = 7. From here we get n - 5 = 7, n = 12.